OpenStudy (zzelinski):

The Coffee Counter charges $11.00 per pound for Kenyan French Roast coffee and $13.00 per pound for Sumatran coffee. How much of each blend should be used to make a 14 pound blend that sells for $12.00 per pound? System of equations, I don't do well with word problems?

OpenStudy (anonymous):

11(K) + 13(S) = (14)(12) where (14)(12) = dollars K + S = 14 total pounds So, just solve those 2 simultaneous equations for K and S.

OpenStudy (anonymous):

All good now, @zzelinski ?

OpenStudy (zzelinski):

Yes, thanks for your help!

OpenStudy (zzelinski):

Haha you're welcome

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @zzelinski

OpenStudy (zzelinski):

Thanks, and no problem :D